OpenStudy (anonymous):

Power (P) is measured in watts using the following formula, where w is the work in joules and t is time in seconds. P - w/t If 5 watts of power are created over a period of 16 seconds, how much work did it take to create the power? A- 11 joules B- 3 joules C- 21 joules D- 80 joules

OpenStudy (anonymous):

substitute p for 5 and then substitute t for 16 5 = w/16 multiply both side by 16 and then you get your answer

OpenStudy (anonymous):

And that would be? sorry i have to hurry i have 10 minutes left on this quiz and still have like 6 questions

OpenStudy (anonymous):

oh well it would be d

OpenStudy (anonymous):

and if you have any more i will just give you a direct answer so you can finish the quiz

OpenStudy (anonymous):

okay thanks so much one second its loading im in a hurry haha

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[m = \frac{ y2 - y1 }{ x2 - x1}\] Solve for y1

OpenStudy (anonymous):

does it give you numbers to substitute

OpenStudy (anonymous):

this is a type your own answer in

OpenStudy (anonymous):

and no ma'am

OpenStudy (anonymous):

i literally have no idea

OpenStudy (anonymous):

im sorry

OpenStudy (anonymous):

okay i got it next one uh solve for L P = 2L + 2w

OpenStudy (anonymous):

A- l = 2p - w B- l = 2/p - w C- l = w/2 - p D- l = p/2 - w

OpenStudy (anonymous):

p-2w over 2 = L the answer might be L = p-2w over 2 but its the same thing

OpenStudy (anonymous):

oh i think it would be c

OpenStudy (anonymous):

says its wrong

OpenStudy (anonymous):

I think its D

OpenStudy (anonymous):

um i think its a i really dont know

OpenStudy (anonymous):

The surface area of a cylinder is given by the formula SA = 2*pi*r^2 + 2*pi*rh. A cylinder has a radius of 8 cm and a surface area of 736 pi cm^2 . Find the height of the cylinder. A-47 cm B-23 cm C-38 cm D-40 cm

OpenStudy (anonymous):

you were right haha

OpenStudy (anonymous):

i have no idea sorry